Jan's age is 3 years less than 2 twice Tritt's age. The sum of their ages is 30. Find their ages.
Answer by macston(5194)

.
T=Tritt's age; J=Jan's age=2T-3
T+J=30
T+(2T-3)=30
3T=33
T=11 ANSWER 1: Tritt is 11 years old.
.
J=2T-3
J=2(11)-3
J=22-3
J=19
ANSWER 2: Jan is 19 years old.